Q4. Two vectors, A and B, are added by means of vector addition to give a resultant vector R R A+ B. The magnitudes of Ä and B are 2.0 m and 7.0 m, respectively; and they can have any orientation. What are the maximum and minimum possible values for the magnitude of R?
Answer #1

Here ,

as for the maximum magnitude of R vector

for maximum value of A + B

when A and B are in same direction

maximum value of A + B = 2 + 7

maximum value of A + B = 9 m

------------------

NOw, for the minimum value of A + B

when A and B are in oposite direction

minimum value of A + B = 7 -2

minimum value of A + B = 5 m

the minimum value of A + B is 5 m